OpenStudy (anonymous):

5 ^3sqrt(9y^2)*^3sqrt(24y)

OpenStudy (anonymous):

please show all work.

OpenStudy (loser66):

hey, it's ambiguous. re-write it in neat if you want to get help

OpenStudy (loser66):

especially at *^.... what is it?

OpenStudy (anonymous):

\[5 \sqrt[3]{9y} \times \sqrt[3]{24y}\] Is this what your problem looks like?

OpenStudy (loser66):

he ran after posting the problem !!???